OpenStudy (anonymous):

Find the equation in slope intercept form passing through the points (7.2, 3.8) and (9.1, 9.1

OpenStudy (anonymous):

HI, FIRST FIND YOUR SLOPE m=(y2-y1)/(x2-x1)=(9.1-3.8)/(9.1-7.2)=5.3/1.9=2.789 now at point(7.2,3.8) use the formula y-y1=m(x-x1) y-3.8=2.789(x-7.2) y=2.789(x-7.2)+3.8 y=2.789x-20.08+3.8 y=2.789x-23.88 ans....